OpenStudy (anonymous):

consider the polynomial function f(x)= x^5+21x^3-100x. determine the zeros of the function by factoring.

OpenStudy (anonymous):

because all the terms have an x you can factor that out first: \[x^5+21x^3-100x = x(x^4+21x^2-100)\] Then think of two numbers that multiply to -100 but add to 21 (which would be 25 and -4): \[x(x^4+21x^2-100) = x(x^2+25)(x^2-4)\] Then factor using a difference of 2 squares: \[x(x^2+25)(x-2)(x+2)\]